OpenStudy (anonymous):

Dr. Wang is jogging down the sidewalk at a speed of 4.20m/s. Her husband realizes that she forgot her lunch bagels, so he runs to the window of their house, which is a height of 35.0m above the street level and directly above the sidewalk, to throw them to her. He throws them horizontally at a time 7.50s after Dr wang has passed below the window, and she catches them on the run. You can ignore air resistance. W/what initial speed must Dr Wang's husband throw the bagels so she can catch them before they hit the ground? Where is Dr Wang when she catches the bagels?

OpenStudy (anonymous):

Let Dr. Wang and her lunch bagels meets at a common point after time t. Distance travelled =4.2(7.50+t) In the same time vertical distance travelled=35 m t =sqrt(2*35/9.8)=0.85 Sec Distance travelled by Dr. Wang=4.2(7.50+0.85)= 35.05m